OpenStudy (anonymous):

Find the slope of the line passing through the given points, when possible. (-2, 5) and (-7, 6)

OpenStudy (anonymous):

slope of the line passing through the points (x1,y1) and (x2,y2) \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

that isn't the answer right @niksva

OpenStudy (anonymous):

what is the slope u r getting after using this concept?

OpenStudy (anonymous):

I haven't got any

OpenStudy (anonymous):

then try to use the concept first

OpenStudy (anonymous):

I am stuck on these slope problems I have no idea how to do them

OpenStudy (anonymous):

\[\frac{ 6-5 }{-7-(-)2 }\] hope u can do now

OpenStudy (anonymous):

@happytales

OpenStudy (anonymous):

Alright he just broke it down for you. Just solve from there.

OpenStudy (anonymous):

but that's what I don't get how do I solve it from there ?

OpenStudy (anonymous):

Just solve. 1/-5 m=-1/5